OpenStudy (anonymous):

7(2x + 6) = 1 i got -41 /14 but not correct

OpenStudy (anonymous):

you are correct

OpenStudy (anonymous):

14x+42=1 14x=1-42 14x=-41 x=-41/14

OpenStudy (anonymous):

ur absolutely ryt thumbs up